# Diffusion Posterior Proximal Sampling for Image Restoration

### v1

@@
Abstract
Diffusion models have demonstrated remarkable efficacy in generating high-quality samples. Existing diffusion-based image restoration algorithms exploit pre-trained diffusion models to leverage data priors, yet they still preserve elements inherited from the unconditional generation paradigm. These strategies initiate the denoising process with pure white noise and incorporate random noise at each generative step, leading to over-smoothed results. In this paper, we introduce a refined paradigm for diffusion-based image restoration. Specifically, we opt for a sample consistent with the measurement identity at each generative step, exploiting the sampling selection as an avenue for output stability and enhancement. Besides, we start the restoration process with an initialization combined with the measurement signal, providing supplementary information to better align the generative process. Extensive experimental results and analyses validate the effectiveness of our proposed approach across diverse image restoration tasks.
Diffusion Model, Image Restoration, Inverse Problem
1 Introduction
@@
Figure 1: Examples of the super-resolution ($\times$4) task to illustrate the efficiency of our method, where $n$ denotes the number of candidate samples, and DPS also refers to the case where $n=1$. Additional results are presented in Appendix F.
Figure 2: High-level illustration of our proposed DPPS ($n$=2). Existing methods employ random sampling from the predicted distribution, our approach takes multiple samples and selecting the one with better data consistency (denoted by $\checkmark$) at each step.
Diffusion models (Vincent, 2011; Ho et al., 2020; Song et al., 2021b) have gained broad attention due to the powerful ability to model complex data distribution, and have been applied to a wide range of tasks such as image generation (Dhariwal & Nichol, 2021; Meng et al., 2021; Rombach et al., 2022), molecule generation (Corso et al., 2023), and natural language processing (Li et al., 2022). Recent research has demonstrated that pre-trained unconditional diffusion models can be effectively employed to address image restoration problems (Song et al., 2022; Kawar et al., 2022; Chung et al., 2022; Wang et al., 2023; Song et al., 2023b) to leverage rich priors in a plug-and-play fashion, and achieve significant advancements.
However, for diffusion-based image restoration algorithms, they still retain a process inherited from unconditional generation. In particular, (i) these methods start generating the image that corresponds to the measurement with a white noise as initialization, and (ii) a fully stochastic noise is introduced at each step of the generation process. We argue that this paradigm is inappropriate for solving image restoration tasks. Firstly, randomness enriches the diversity of the generated samples in unconditional generation (Liu et al., 2020; Xiao et al., 2021). However, for image restoration tasks where the identity information of the measurement needs to be preserved, randomness leads to uncontrollable generation outcomes (Cao et al., 2023). Secondly, in the generative process of diffusion-based model, the current state is randomly sampled (Chung et al., 2023a; Song et al., 2023b) from the predicted distribution without any correction. Considering the Markovian reverse process, the sampling quality of the next state heavily depends on the current sampling result. If the current result deviates considerably from expectations, the subsequent generative process will encounter a significant exposure bias problem (Ning et al., 2023; Li et al., 2023). As a consequence, randomness introduces fluctuations and instability into the generative process, ultimately leading to over-smoothed results.
@@
  • We introduce an efficient proximal sampling strategy that aligns with measurement identity to solve diffusion-based image restoration problems.
  • Extensive experiments are conducted to demonstrate our superior performance in image restoration tasks, and abundant ablation analysis verifies the validity of the proposed method.
2 Background
2.1 Denoising Diffusion Probabilistic Models
@@
$q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbf{I}),$ (1)
where $\mathcal{N}$ denotes the Gaussian distribution, and $\beta_{t}$ is a pre-defined parameter increasing with $t$. Through reparameterization, $\mathbf{x}_{t}$ satisfies the following conditional distribution $q(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I})$, where $\bar{\alpha}_{t}=\prod_{i=0}^{t}\alpha_{i}$ and $\alpha_{t}=1-\beta_{t}$.
@@
$p_{\mathbf{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t-1};\bm{\mu}_{\mathbf{\theta}}(\mathbf{x}_{t},t),\sigma_{t}\mathbf{I}).$ (2)
@@
$\mathbf{L}({\mathbf{\theta}})=\mathbb{E}_{\mathbf{x}_{0}\sim q(\mathbf{x}_{0}),{\bm{\epsilon}}\sim{\cal N}(\mathbf{0},\mathbf{I}),t\sim\mathbb{U}(\{1,...,T\})}[\|{\bm{\epsilon}}-{\bm{\epsilon}}_{\mathbf{\theta}}(\mathbf{x}_{t},t)\|^{2}].$ (3)
where $\mathbf{x}_{0}$ is sampled from training data and $\mathbf{x}_{t}\sim q(\mathbf{x}_{t}|\mathbf{x}_{0})$. Once we have access to the well-trained $\bm{\epsilon}_{\mathbf{\theta}}(\mathbf{x}_{t},t)$, a clean sample can be derived by evaluating the generative reverse process Eq. (2) step by step.
@@
$\mathbf{L}({\mathbf{\theta}})=\mathbb{E}_{\mathbf{x}_{0},{\bm{\epsilon}},t}[||\mathbf{s}_{{\mathbf{\theta}}}(\mathbf{x}_{t},t)-\nabla_{\mathbf{x}_{t}}\log q(\mathbf{x}_{t}|\mathbf{x}_{0})||^{2}].$ (4)
Algorithm 1 DPS
0:  $T$, $\mathbf{y}$, $\lambda_{t}\}_{t=1}^{T}$, ${\{\sigma_{t}\}_{t=1}^{T}}$
1:  $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
2:  for $t=T$ to $1$ do
3:     compute $\bm{\mu}_{\mathbf{\theta}}(\mathbf{x}_{t},t,\mathbf{y})$ via (8)
4:     if $t>1$ then ${\mathbf{z}}_{t}\sim{\mathcal{N}}(\mathbf{0},\mathbf{I})$, else $\mathbf{z}_{t}=\mathbf{0}$
5:     $\mathbf{x}_{t-1}\leftarrow\bm{\mu}_{\mathbf{\theta}}(\mathbf{x}_{t},t,\mathbf{y})+\sigma_{t}{\mathbf{z}}_{t}$
6:  end for
7:  return $\hat{\mathbf{x}}_{0}$
2.2 Diffusion Based Solvers for Image Restoration
@@
$\mathbf{y}=\mathcal{A}(\mathbf{x}_{0})+\mathbf{n},\quad\mathbf{y},\mathbf{n}\in{\mathbb{R}}^{n},\,\mathbf{x}\in{\mathbb{R}}^{d},$ (5)
@@
$\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}|\mathbf{y})=\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})+\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}|\mathbf{x}_{t}).$ (6)
The first prior term can be approximated by a well-trained score network. However, the analytic solution of the second likelihood term is computationally intractable, since there only exists an explicit connection between $\mathbf{y}$ and $\mathbf{x}_{0}$. To solve this dilemma, Chung et al. (2023a) propose diffusion posterior sampling (DPS) to approximate the likelihood using a Laplacian approximation: $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}|\mathbf{x}_{t})\simeq\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}|\hat{\mathbf{x}}_{0|t}),$ where $\hat{\mathbf{x}}_{0|t}$ is the denoised estimate via Tweedie’s formula (Stein, 1981; Efron, 2011). Consequently, the generative distribution $p_{\mathbf{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{y})$ can be modeled as:
$p_{\mathbf{\theta}}(\mathbf{x}_{t-1}|{\mathbf{x}}_{t},\mathbf{y}):=\mathcal{N}\big{(}\mathbf{x}_{t};~{}{\bm{\mu}}_{\mathbf{\theta}}(\mathbf{x}_{t},t,\mathbf{y}),~{}{\sigma_{t}}\mathbf{I}\big{)}.$ (7)
The mean of Gaussian distribution ${\bm{\mu}}_{\mathbf{\theta}}(\mathbf{x}_{t},t,\mathbf{y})$ is obtained by:
${\bm{\mu}}_{\mathbf{\theta}}(\mathbf{x}_{t},t,\mathbf{y})=$ $\frac{1}{\sqrt{\alpha_{t}}}\Big{(}\mathbf{x}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}{\bm{\epsilon}}_{\mathbf{\theta}}(\mathbf{x}_{t},t)\Big{)}$ (8)
@@
where $\hat{\mathbf{x}}_{0|t}=\frac{1}{\sqrt{\bar{\alpha}_{t}}}\big{(}\mathbf{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}_{\mathbf{\theta}}(\mathbf{x}_{t},t)\big{)}$, and $\lambda_{t}$ is a tunable step size. We detail the methodology  [^1] There is a slight difference between our implementation and the original DPS, but the experimental performance shows little difference. in Algorithm 1.
2.3 Random Sampling Induces Uncertainties and Error Accumulation
@@
Existing approaches address image restoration or inverse problems following a process derived from unconditional generation. Specifically, the process initiates the denoising process with pure white noise and incorporates random noise at each reverse step. We contend that randomness is unsuitable for restoration problems that demand the preservation of measurement identities, such as super-resolution or deblurring. Furthermore, the cumulative impact of random noise at each step results in the smoothing of output, thereby yielding low-quality generated samples. Recent investigations (Ma et al., 2023; Everaert et al., 2023) align with our findings, highlighting that randomness introduces instability and fluctuations, ultimately culminating in suboptimal samples.
@@
Besides, we observe a discrepancy in the inputs provided to the noise prediction network $\bm{\epsilon}_{\mathbf{\theta}}$. During the training phase, the network is fed with ground truth training samples. However, in the inference stage, the input $\mathbf{x}_{t}$ is randomly sampled from the predicted distribution. In cases where the predicted distribution is inaccurate, random sampling can exacerbate the deviation from the expected values, introducing substantial exposure bias (Ning et al., 2023) to the generative process. While the data consistency update $\nabla_{\mathbf{x}_{t}}||\mathbf{y}-\mathcal{A}\hat{\mathbf{x}}_{0|t}||$ does offer mitigation by adjusting the sample to align with the measurement, it demands delicate design and specifically-tuned parameters (Song et al., 2023c). And the optimal parameter values vary across datasets and tasks. Consequently, the selection of sampling choices holds significance as it directly influences the input to the network, thereby affecting the generation quality.
3 Diffusion Posterior Proximal Sampling
3.1 Proximal Sampling at Each Step

### v2

@@
Abstract.
Diffusion models have demonstrated remarkable efficacy in generating high-quality samples. Existing diffusion-based image restoration algorithms exploit pre-trained diffusion models to leverage data priors, yet they still preserve elements inherited from the unconditional generation paradigm. These strategies initiate the denoising process with pure white noise and incorporate random noise at each generative step, leading to over-smoothed results. In this paper, we present a refined paradigm for diffusion-based image restoration. Specifically, we opt for a sample consistent with the measurement identity at each generative step, exploiting the sampling selection as an avenue for output stability and enhancement. The number of candidate samples used for selection is adaptively determined based on the signal-to-noise ratio of the timestep. Additionally, we start the restoration process with an initialization combined with the measurement signal, providing supplementary information to better align the generative process. Extensive experimental results and analyses validate that our proposed method significantly enhances image restoration performance while consuming negligible additional computational resources.
Diffusion Model
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3664647.3681556isbn: 979-8-4007-0686-8/24/10ccs: Computing methodologies Computer vision problemsccs: Computing methodologies Reconstructionccs: Computing methodologies Machine learning
1. Introduction
@@
Figure 1. Examples of the super-resolution ($\times$4) task to illustrate the efficiency of our method, where $n$ denotes the number of candidate samples, and DPS also refers to the case where $n=1$.
[…]
Examples of the super-resolution ($\times$4) task to illustrate the efficiency of our method, where $n$ denotes the number of candidate samples, and DPS also refers to the case where $n=1$.
Figure 2. High-level illustration of our proposed DPPS ($n$=2). Existing methods employ random sampling from the predicted distribution, our approach takes multiple samples and selecting the one with better data consistency (denoted by $\checkmark$) at each step.
[…]
High-level illustration of our proposed DPPS ($n$=2). Existing methods employ random sampling from the predicted distribution, our approach takes multiple samples and selecting the one with better data consistency (denoted by $\checkmark$) at each step.
Diffusion models (Vincent, 2011; Ho et al., 2020; Song et al., 2021b) have gained broad attention due to the powerful ability to model complex data distribution, and have been applied to a wide range of tasks such as image generation (Dhariwal and Nichol, 2021; Meng et al., 2021; Rombach et al., 2022), molecule generation (Corso et al., 2023), and natural language processing (Li et al., 2022). Recent research has demonstrated that pre-trained unconditional diffusion models can be effectively employed to address image restoration problems (Song et al., 2022; Kawar et al., 2022; Chung et al., 2022; Wang et al., 2023; Song et al., 2023b) to leverage rich priors in a plug-and-play fashion, and achieve significant advancements.
However, for diffusion-based image restoration algorithms, they still retain a process inherited from unconditional generation. In particular, (i) these methods start generating the image that corresponds to the measurement with a white noise as initialization, and (ii) a fully stochastic noise is introduced at each step of the generation process. We argue that this paradigm is inappropriate for solving image restoration tasks. Firstly, randomness enriches the diversity of the generated samples in unconditional generation (Liu et al., 2020; Xiao et al., 2021). However, for image restoration tasks where the identity information of the measurement needs to be preserved (Yao et al., 2024), randomness leads to uncontrollable generation outcomes (Cao et al., 2023). Secondly, in the generative process of diffusion-based model, the current state is randomly sampled (Chung et al., 2023a; Song et al., 2023b) from the predicted distribution without any correction. Considering the Markovian reverse process, the sampling quality of the next state heavily depends on the current sampling result. If the current result deviates considerably from expectations, the subsequent generative process will encounter a significant exposure bias problem (Ning et al., 2023b; Li et al., 2023). As a consequence, randomness introduces fluctuations and instability into the generative process, ultimately leading to over-smoothed results.
@@
  • We propose an efficient proximal sampling strategy that aligns with measurement identity to solve diffusion-based image restoration problems.
  • Extensive experimental results demonstrate that our proposed method outperforms other state-of-the-art methods in image restoration performance, requiring only minimal additional computation.
2. Background
2.1. Denoising Diffusion Probabilistic Models
@@
(1) $q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbf{I}),$
where $\mathcal{N}$ denotes the Gaussian distribution, and $\beta_{t}$ is a pre-defined parameter increasing with $t$. Through reparameterization, $\mathbf{x}_{t}$ satisfies the following conditional distribution $q(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},\\ (1-\bar{\alpha}_{t})\mathbf{I})$, where $\bar{\alpha}_{t}=\prod_{i=0}^{t}\alpha_{i}$ and $\alpha_{t}=1-\beta_{t}$.
@@
(2) $p_{\mathbf{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t-1};\boldsymbol{\mu}_{\mathbf{\theta}}(\mathbf{x}_{t},t),\sigma_{t}^{2}\mathbf{I}).$
@@
(3) $\mathbf{L}({\mathbf{\theta}})=\mathbb{E}_{\mathbf{x}_{0}\sim q(\mathbf{x}_{0}),{\boldsymbol{\epsilon}}\sim{\mathcal{N}}(\mathbf{0},\mathbf{I}),t\sim\mathbb{U}(\{1,...,T\})}[\|{\boldsymbol{\epsilon}}-{\boldsymbol{\epsilon}}_{\mathbf{\theta}}(\mathbf{x}_{t},t)\|^{2}].$
where $\mathbf{x}_{0}$ is sampled from training data and $\mathbf{x}_{t}\sim q(\mathbf{x}_{t}|\mathbf{x}_{0})$. Once we have access to the well-trained $\boldsymbol{\epsilon}_{\mathbf{\theta}}(\mathbf{x}_{t},t)$, a clean sample can be derived by evaluating the generative reverse process Eq. (2) step by step.
@@
(4) $\mathbf{L}({\mathbf{\theta}})=\mathbb{E}_{\mathbf{x}_{0},{\boldsymbol{\epsilon}},t}[||\mathbf{s}_{{\mathbf{\theta}}}(\mathbf{x}_{t},t)-\nabla_{\mathbf{x}_{t}}\log q(\mathbf{x}_{t}|\mathbf{x}_{0})||^{2}].$
2.2. Diffusion Based Solvers for Image Restoration
@@
(5) $\mathbf{y}=\mathcal{A}(\mathbf{x}_{0})+\mathbf{n},\quad\mathbf{y},\mathbf{n}\in{\mathbb{R}}^{n},\,\mathbf{x}\in{\mathbb{R}}^{d},$
@@
(6) $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}|\mathbf{y})=\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})+\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}|\mathbf{x}_{t}).$
The first prior term can be approximated by a well-trained score network. However, the analytic solution of the second likelihood term is computationally intractable, since there only exists an explicit connection between $\mathbf{y}$ and $\mathbf{x}_{0}$. To solve this dilemma, Chung et al. (2023a) propose diffusion posterior sampling (DPS) to approximate the likelihood using a Laplacian approximation: $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}|\mathbf{x}_{t})\simeq\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}|\hat{\mathbf{x}}_{0|t}),$ where $\hat{\mathbf{x}}_{0|t}$ is the denoised estimate via Tweedie’s formula (Stein, 1981; Efron, 2011). Consequently, the generative distribution $p_{\mathbf{\theta}}(\mathbf{x}_{t-1}|\\ \mathbf{x}_{t},\mathbf{y})$ can be modeled as:
(7) $p_{\mathbf{\theta}}(\mathbf{x}_{t-1}|{\mathbf{x}}_{t},\mathbf{y}):=\mathcal{N}\big{(}\mathbf{x}_{t};~{}{\boldsymbol{\mu}}_{\mathbf{\theta}}(\mathbf{x}_{t},t,\mathbf{y}),~{}{\sigma_{t}^{2}}\mathbf{I}\big{)}.$
The mean of Gaussian distribution ${\boldsymbol{\mu}}_{\mathbf{\theta}}(\mathbf{x}_{t},t,\mathbf{y})$ is obtained by:
(8) ${\boldsymbol{\mu}}_{\mathbf{\theta}}(\mathbf{x}_{t},t,\mathbf{y})=$ $\frac{1}{\sqrt{\alpha_{t}}}\Big{(}\mathbf{x}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}{\boldsymbol{\epsilon}}_{\mathbf{\theta}}(\mathbf{x}_{t},t)\Big{)}$
@@
where $\hat{\mathbf{x}}_{0|t}=\frac{1}{\sqrt{\bar{\alpha}_{t}}}\big{(}\mathbf{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}{\boldsymbol{\epsilon}}_{\mathbf{\theta}}(\mathbf{x}_{t},t)\big{)}$, and $\lambda_{t}$ is a tunable step size.
3. Diffusion Posterior Proximal Sampling
3.1. Random Sampling Induces Uncertainties and Error Accumulation
@@
Existing approaches (Song et al., 2022; Kawar et al., 2022; Chung et al., 2022; Wang et al., 2023; Song et al., 2023b) address image restoration or inverse problems following a process derived from unconditional generation. Specifically, the process (i) initiates the denoising process with pure white noise, and (ii) incorporates random noise at each reverse step. We contend that randomness is unsuitable for restoration problems that demand the preservation of measurement identities, such as super-resolution or deblurring. Furthermore, the cumulative impact of random noise at each step results in the smoothing of output, thereby yielding low-quality generated samples. Recent investigations (Ma et al., 2023; Everaert et al., 2023) align with our findings, highlighting that randomness introduces instability and fluctuations, ultimately culminating in suboptimal samples.
@@
On the other hand, we observe a discrepancy in the inputs provided to the noise prediction network $\boldsymbol{\epsilon}_{\mathbf{\theta}}$. During the training phase, the network is fed with ground truth training samples. However, in the inference stage, the input $\mathbf{x}_{t}$ is randomly sampled from the predicted distribution. In cases where the predicted distribution is inaccurate, random sampling can exacerbate the deviation from the expected values, introducing substantial exposure bias (Ning et al., 2023b) to the generative process. While the data consistency update $\nabla_{\mathbf{x}_{t}}||\mathbf{y}-\mathcal{A}\hat{\mathbf{x}}_{0|t}||$ does offer mitigation by adjusting the sample to align with the measurement, it demands delicate design and specifically-tuned parameters (Song et al., 2023c). And the optimal parameter values vary across datasets and tasks. Consequently, the selection of sampling choices holds significance as it directly influences the input to the network, thereby affecting the generation quality.
3.2. Proximal Sampling at Each Step
To tackle the challenges posed by random sampling, in this paper, we propose to extract multiple candidate samples from the predicted distribution, and select the most proximal one (Parikh and Boyd, 2014) to our anticipated target. Our motivation comes from the following idea: considering $\mathbf{x}_{0}$ is an ideal but unknown solution for the image restoration problem, the sample taken from the predicted distribution $\mathbf{x}_{t-1}\sim p_{\mathbf{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{y})$ should be close to posterior $q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})$, as it models the desired reverse process.
The mathematical formulation of our selection process is given by:
(9) $\mathbf{x}_{t-1}=$ $\operatorname*{arg\,min}_{\mathbf{x}_{t-1}^{i}}\|\mathbf{x}_{t-1}^{i}-\mathbf{x}_{t-1}^{*}\|_{2}^{2}$
where $\mathbf{x}_{t-1}^{i}\sim p_{\mathbf{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{y}),~{}i\in[0,n-1]$, and $\mathbf{x}_{t-1}^{*}$ is our proposed deterministic solution 111the symbol means an ideal but unknown solution. via DDIM (Song et al., 2021a) with unknown $\mathbf{x}_{0}$ (see supplementary material for detailed derivation):
(10) $\mathbf{x}_{t-1}^{*}=$ $\sqrt{\bar{\alpha}_{t-1}}{\mathbf{x}}_{0}+\sqrt{1-\bar{\alpha}_{t-1}}\cdot\frac{\mathbf{x}_{t}-\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}}{\sqrt{1-\bar{\alpha}_{t}}}$
$=$ $C_{1}\mathbf{x}_{t}+C_{2}\mathbf{x}_{0}.$
Here $C_{1}={\sqrt{1-\bar{\alpha}_{t-1}}}/{\sqrt{1-\bar{\alpha}_{t}}}$ and $C_{2}=\sqrt{\bar{\alpha}_{t-1}}-{\sqrt{\bar{\alpha}_{t}}\sqrt{1-\bar{\alpha}_{t-1}}}/\\ {\sqrt{1-\bar{\alpha}_{t}}}$ are introduced for simplicity.
Figure 3. Conceptual illustration of our DPPS. Our method extract multiple candidate samples from the predicted distribution and choose the one with the highest measurement consistency.
[…]
Conceptual illustration of our DPPS. Our method extract multiple candidate samples from the predicted distribution and choose the one with the highest measurement consistency.
However, accessing $\mathbf{x}_{0}$ during inference is infeasible, rendering $\mathbf{x}_{t-1}^{*}$ also unknown. Now, one important contribution of this paper is to project Eq. (9) onto the measurement subspace. This is feasible as $\mathcal{A}\mathbf{x}_{t-1}^{*}$ can be approximated under the condition $\mathbf{y}\approx\mathcal{A}\mathbf{x}_{0}$ when the degradation operator is linear and $\sigma_{y}$ is assumed within a moderate range.
(11) $\mathcal{A}\mathbf{x}_{t-1}^{*}\approx C_{1}\mathcal{A}\mathbf{x}_{t}+C_{2}\mathbf{y}.$
(12) $\mathbf{x}_{t-1}=\operatorname*{arg\,min}_{\mathbf{x}_{t-1}^{i}}\|\mathcal{A}(\mathbf{x}_{t-1}^{i}-\mathbf{x}_{t-1}^{*})\|_{2}^{2}$
Since $\mathbf{x}_{0}$ is expected to be more accurate than the denoised result $\hat{\mathbf{x}}_{0|t}$, the measurement $\mathbf{y}\approx\mathcal{A}\mathbf{x}_{0}$ can provide a stable and strong supervisory signal to correct the sampling result. Then by means of sample selection, we effectively control the result $\mathbf{x}_{t-1}$ within a more desirable region, even though we have no access to the $\mathbf{x}_{0}$.
As $\mathcal{A}$ is irreversible and complex in numerous scenarios (Chen et al., 2021, 2023), obtaining its pseudo-inverse poses challenges. Therefore, we propose to randomly sample $n$ candidates $\mathbf{x}_{t-1}^{i},i\in[0,n-1]$ from the predicted distribution $p_{\mathbf{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{y})$, and choose the one with the minimal deviation from our anticipated values. Finally, Eq. (9) becomes:
(13) $\mathbf{z}_{t}=$ $\operatorname*{arg\,min}_{\mathbf{z}_{t}^{i}}\|\mathcal{A}\left(\boldsymbol{\mu}_{\mathbf{\theta}}(\mathbf{x}_{t},t,\mathbf{y})+\sigma_{t}\mathbf{z}_{t}^{i}-C_{1}\mathbf{x}_{t}\right)-C_{2}\mathbf{y}\|_{2}^{2}.$
The process of sample selection is also denoted as noise selection, given by $\mathbf{x}_{t-1}^{i}=\boldsymbol{\mu}_{\mathbf{\theta}}(\mathbf{x}_{t},t,\mathbf{y})+\sigma_{t}\mathbf{z}_{t}^{i}$. We detail the full algorithm in Algorithm 1, and name our algorithm Diffusion Posterior Proximal Sampling (DPPS). The chosen $\mathbf{x}_{t-1}$ is referred to as the proximal sample due to the similarity to the proximal operator (Parikh and Boyd, 2014).
3.3. Adaptive Sampling Frequency
It is commonly perceived that the generative process of diffusion does not exhibit uniform significance across all timesteps (Karras et al., 2022; Kingma and Gao, 2023; Esser et al., 2024). Based on this insight, we adaptively adjust the number of candidate samples during the generative process according to the signal-to-noise ratio to enhance image quality and reduce computational cost. Through extensive experimentation, we discovered that using a larger number of candidate samples in the final stage of generation yields better results. The number of candidate samples $n$ at each step can be expressed as:
(14) $n=\max(\lfloor N_{max}*(1-e^{-\lambda_{t}})\rfloor,~{}2),$
where $t\in[0,999]$ is the timestep, $\lambda_{t}=\frac{\bar{\alpha}_{t}}{1-\bar{\alpha}_{t}}$ represents the signal-to-noise ratio, and $N_{max}$ is a hyper-parameter that adjusts the maximum sampling frequency. The minimum frequency is set to $2$, ensuring our proximal sampling strategy.
3.4. Aligned Initialization
Recent studies have pointed out that initializition have a significant impact on the generated results (Singh et al., 2022; Cao et al., 2023; Ma et al., 2023), and the discrepancy between the two distributions account for a discretization error (Benton et al., 2023). In this paper, inspired by (Everaert et al., 2023), we simply initialize the sample in the same way as during training, making the best use of the available measurement
(15) $\mathbf{x}_{T}=\sqrt{\bar{\alpha}_{T}}\mathcal{A}^{T}\mathbf{y}+\sqrt{1-\bar{\alpha}_{T}}\boldsymbol{\epsilon}\quad\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}).$